November 1996  ICTCM IX (Reno)  15

Interpretation of Solution

• terminal velocities: – free-fall: v_{ff} ≈ 49 m/s

– final descent: v_{impact} ≈ 7m/s • position:

– deployment at 1300 m

– land after 241 s

• acceleration:

– jump discontinuity at t_{d}

– magnitude: 6 G (all values approximate)